OpenStudy (anonymous):

integral of sqrt(1+cos2x)dx

OpenStudy (anonymous):

\[\int\limits \sqrt{\left( 1+\cos 2x \right)}dx=\int\limits \sqrt{2\cos ^2x}~dx=\sqrt{2}\int\limits \cos x~dx+c\] =?

OpenStudy (anonymous):

@surjithayer where did the one go then when you substituted the 2cos^2x in for cos2x?

terenzreignz (terenzreignz):

It's an identity, @professional_loser_ ^_^ \[\Large 1 = \cos^2(x) + \sin^2(x)\]\[\Large \cos(2x) = \cos^2(x) - \sin^2(x)\] And just add up these two equations :D

OpenStudy (anonymous):

thank you! I completely blanked on that identity, lol. @terenzreignz